# Characterizing the Distinguishability of Product Distributions through Multicalibration

### v3

@@ 1.1 Main Results
    $X_{0}^{\otimes k}$ and $X_{1}^{\otimes k}$ are $(d_{TV}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-2k\cdot\varepsilon)$-distinguishable by circuits of size $s^{\prime}=O(sk/\varepsilon^{2})+\text{poly}(k/\varepsilon)$.
@@
    The statements above also hold with $\widetilde{X}_{1}=X_{1}$, but $s^{\prime}=O(sk/\varepsilon^{4})+\text{poly}(k/\varepsilon)$.
@@
Item 2 follows directly from Item 1 via hybrid argument, so the value of Theorem 1.1 is that it provides not only an upper bound on (in)distinguishability but a matching lower bound, via Item 3. Thinking of $s=n^{\omega(1)},k\leq\mathrm{poly}(n),\varepsilon=1/s$, we see that up to a negligible additive term $(k\varepsilon)$, and a polynomial change in circuit size ($s^{\prime}$ vs. $s$), the computational indistinguishability of $X_{0},X_{1}$ under multiple samples is tightly captured by the information-theoretic indistinguishability of $\widetilde{X}_{0},\widetilde{X}_{1}$ under multiple samples. In particular, $k=\Theta\left(1/d^{2}_{\mathrm{H}}(\widetilde{X}_{0},\widetilde{X}_{1})\right)$ samples are both necessary and sufficient to distinguish $X_{0},X_{1}$ with constant advantage. We can abstract this latter corollary as follows:
@@
    $X_{0}^{\otimes k},X_{1}^{\otimes k}$ are $(1-e^{-k\delta^{2}}-2k\varepsilon)$ distinguishable by circuits of size $s^{\prime}=O(sk/\varepsilon^{2})+\mathrm{poly}(k/\varepsilon)$.
Thus, the number of samples needed to distinguish $X_{0},X_{1}$ with constant advantage is $\Theta(1/\delta^{2})$, where $\delta$ is the pseudo-Hellinger distance. As an immeidate consequence, just like the traditional notion of Hellinger distance in the statistical distinguishability setting, our notion of pseudo-Hellinger distance gives an instance-optimal characterization of the computational distinguishability of any pairs of random variables. This is a key benefit of our work over prior works [HR08, Gei22].
@@
    $X_{0}^{\otimes k},\mathcal{U}^{\otimes k}$ are $(1-e^{-\Omega(k\min\{\delta,1\})}-2k\varepsilon)$ distinguishable by circuits of size $s^{\prime}=O(sk/\varepsilon^{4})+\mathrm{poly}(k/\varepsilon)$.
@@ 1.2 Asymptotic Complexity Formulations
We say that ensembles $X,Y$ have computational sample complexity at least $k:\mathbb{N}\rightarrow\mathbb{N}$ for $k(n)=n^{O(1)}$ if $X^{k}\equiv^{\mathrm{comp}}_{1/2}Y^{k}$ (where $X^{k}=\{X_{n}^{k(n)}\}$). We denote this as $X\equiv_{k}^{cS}Y$.
@@
    If $X\equiv_{k}^{cS}Y$, then $X\equiv^{cH}_{\delta}Y$ for some $\delta=O\left(\frac{1}{\sqrt{k(n)}}\right)$.
@@ 1.3 Technical Overview
Our work builds on the recent work of Casacuberta, Dwork, and Vadhan [CDV24], who showed how the recent notion of multicalibration from the algorithmic fairness literature [HKRR18] leads to simpler and more illuminating proofs of a number of fundamental known results in complexity theory and cryptography, such as the Impagliazzo Hardcore Lemma [Imp95], the Dense Model Theorem [RTTV08], and characterizations of pseudoentropy [VZ12]. We show how multicalibration can be used to derive new results about computational indistinguishability (Theorems 1.1 and 1.2) and in general reduce computational reasoning to information-theoretic reasoning. Specifically, applying the Multicalibration Theorem [HKRR18] to distinguishing problems in a way inspired by [CDV24], we obtain the following. Let $\widetilde{X}_{b|p(\widetilde{X}_{b})=y}$ be the distribution where $x$ is sampled according to $\widetilde{X}_{b}$ conditioned on $p(x)=y$, for some function $p:\mathcal{X}\to[m]$. Let $p(X_{0})$ be the distribution over $[m]$ where for $y\in[m]$, the probability that $p(X_{0})=y$ is exactly $\sum_{x\in p^{-1}(y)}X_{0}(x)$.
@@
    $p$ is computable by circuits of size $O(s/\varepsilon^{2})$.
@@ 1.3.1 Multicalibrated Partitions
We use a partition-based characterization of multicalibration as used in previous works [GKSZ22, GKR+22, CDV24]:
@@
Let $\mathcal{X}$ be a finite domain, for an integer $s\in\mathbb{Z}^{+}$, let $\mathcal{F}^{(s)}$ be the class of functions $f:\mathcal{X}\rightarrow\{0,1\}$ computable by size $s$ circuits, let $g:\mathcal{X}\rightarrow[0,1]$ be an arbitrary function, $\mathcal{D}$ be a probability distribution over $\mathcal{X}$, and $\varepsilon,\gamma>0$. Then, there exists a $(\mathcal{F}_{s},\varepsilon,\gamma)$-approximately multicalibrated partition $\mathcal{P}$ of $\mathcal{X}$ for $g$ on $\mathcal{D}$ such that $\mathcal{P}\in\mathcal{F}^{(W,k)}$, where $W=O(s/\varepsilon^{2})+O(1/(\varepsilon^{4}\gamma)\cdot\log(|\mathcal{X}|/\varepsilon))$.
@@
Recently, this perspective on multicalibration has proven to be quite fruitful, with applications to complexity theory (specifically, towards a better understanding of hardcore sets, the Dense Model Theorem, and pseudoentropy) [CDV24] and machine learning [GKR+22]. Philosophically, the work of [CDV24] showed how to reprove (and strengthen) theorems of average-case hardness and indistinguishability using multicalibration. One consequence of their work was a “loose” intuition that multicalibration translates questions about computational indistinguishability into questions about statistical indistinguishability. Our goal is to show this translation more clearly and generally, and underline directly how multicalibration reduces an understanding of computational indistinguishability to an understanding of statistical indistinguishability (as captured by Theorem 1.1).
@@ 1.3.2 Invoking Multicalibration
With this definition, we can provide an informal overview of the proof of Theorem 1.5. We start with item (b), as its proof is essentially definitional. Recall that this item states that $p(X_{0})$ and $p(\widetilde{X}_{0})$ are identically distributed (and analogously for $X_{1}$). This follows because in the construction of the random variable $\widetilde{X}_{0}$, we sample an element from part $P_{i}$ with probability $\Pr[X_{0}\in P_{i}]$, and thus necessarily, the probability mass placed on each part $P_{i}$ is identical between the two distributions. Next, we can also see that part (c) is definitional. In the construction of $\widetilde{X}_{0}$ and $\widetilde{X}_{1}$, conditioned on being in the piece $P_{i}$, each of the marginal distributions is exactly $\mathcal{D}_{P_{i}}$, and thus the marginal distributions match. Next, to conclude item (d) of Theorem 1.5, recall that the partition function $p$ is exactly returned by the Multicalibration Theorem. Here, we can take advantage of the fact that multicalibrated partitions are efficiently computable given oracle access to the underlying set of test functions (in this case, size $\leq s$ circuits). By replacing these oracles with the size $\leq s$ circuits, this yields an overall complexity of $O(s/\varepsilon^{2})$ for the circuit size required to compute $p$. Finally, it remains to prove item (a) of Theorem 1.5. Because the partition $\mathcal{P}$ that is returned is multicalibrated with respect to our function $g$, by applying Theorem 2.15 of [CDV24] and following the approach of the DMT+++++ + proof of [CDV24], we can show that for every part $P_{i}\in\mathcal{P}$ and every circuit of size $\leq s$, the distributions $X_{0}|_{P_{i}}$ and $\widetilde{X}_{0}|_{P_{i}}$ are close to indistinguishable (and likewise for $X_{1}$). By combining this “local” indistinguishability across all parts of the partition, this then yields the overall indistinguishability between $\widetilde{X}_{0},\widetilde{X}_{1}$.
@@
It is worth comparing our use of the Multicalibration Theorem to the use of Impagliazzo’s Hardcore Theorem [Imp95, Hol05] in some past works on computational indistinguishability and pseudorandomness [HR08, MT10]. Applied to the same function $g$ above, the Hardcore Theorem says that if $X_{0},X_{1}$ are $(s^{\prime},1-\delta)$indistinguishable, for $s^{\prime}=O(s\cdot\mathrm{poly}(1/\varepsilon,1/\delta)$, then there is a subset $H\subseteq X$ such that $\Pr_{x\in X_{0}}[x\in H]=\Pr_{x\in X_{1}}[x\in H]=\delta$ such that $X_{0}|_{H}$ and $X_{1}|_{H}$ are $(s,\varepsilon)$-indistinguishable. Replacing $X_{0}|_{H}$ and $X_{1}|_{H}$ with their average as above, we obtain $\widetilde{X}_{0},\widetilde{X}_{1}$ that are $(s,\varepsilon)$-indistinguishable from $X_{0},X_{1}$ respectively, and such that $d_{\mathrm{TV}}(\widetilde{X}_{0},\widetilde{X}_{1})\leq 1-\delta.$ Thus, the Hardcore Lemma tightly captures the single-sample computational indistinguishability of $X_{0}$ and $X_{1}$ by the single-sample information-theoretic indistinguishability of $\widetilde{X}_{0}$ and $\widetilde{X}_{1}$. But it does not seem to capture enough information to obtain an instance-optimal characterization of the multiple-sample indistinguishability, as we do. Concretely, the Hardcore Lemma requires assuming some initial average-case hardness of $g$ (which comes from the weak indistinguishability of $X_{0}$ and $X_{1}$) and only gives us indistinguishability from an information-theoretic counterpart on a small subset $H\subseteq\mathcal{X}$, which is not guaranteed to be efficiently recognizable. In contrast, the partition given by the Multicalibration Theorem covers the entire domain $\mathcal{X}$ with efficiently recognizable sets, and this is crucial for our instance-optimal characterization.
@@ 2.1 Distinguishability
$d_{H}^{2}\left(X,\mathcal{U}\right)=1-\sqrt{\frac{2^{\mathrm{H}_{1/2}\left(X_{0}\right)}}{|\mathcal{X}|}}.$
@@
$d_{H}^{2}\left(X,\mathcal{U}\right)=\frac{1}{2}\sum_{i\in\mathcal{X}}\left(\sqrt{\mathbb{P}[X_{0}=i]}-\frac{1}{\sqrt{|\mathcal{X}|}}\right)^{2}=\frac{1}{2}\left(1+1-2\sqrt{\frac{\mathbb{P}[X_{0}=i]}{|\mathcal{X}|}}\right)$
$=1-\frac{1}{\sqrt{|\mathcal{X}|}}\sum_{i\in\mathcal{X}}\sqrt{\mathbb{P}[X_{0}=i]}=1-\sqrt{\frac{2^{\mathrm{H}_{1/2}\left(X_{0}\right)}}{|\mathcal{X}|}}.$
@@ 2.2 Multicalibration
Let $\mathcal{X}$ be a finite domain, $\mathcal{F}$ a class of functions, with $f\in\mathcal{F}:f:\mathcal{X}\rightarrow[0,1]$, $g:\mathcal{X}\rightarrow[0,1]$ an arbitrary function, $\mathcal{D}$ a probability distribution over $\mathcal{X}$, and $\varepsilon,\gamma>0$. Then, there exists a $(\mathcal{F},\varepsilon,\gamma)$-approximately multicalibrated partition $\mathcal{P}$ of $\mathcal{X}$ for $g$ on $\mathcal{D}$ such that $\mathcal{P}\in\mathcal{F}_{q,t,k}$ for $t=O(1/(\varepsilon^{4}\gamma)\cdot\log(|\mathcal{X}|/\varepsilon))$, $q=O(1/\varepsilon^{2})$, and $k=O(1/\varepsilon)$.
@@
    $p$ is computable by functions in $\mathcal{F}_{O(1/\varepsilon^{2}),O(1/(\varepsilon^{6})\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))}$.

### v5

@@ 1.1 Main Results
    $X_{0}^{\otimes k}$ and $X_{1}^{\otimes k}$ are $(d_{TV}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-2k\cdot\varepsilon)$-distinguishable by circuits of size $s^{\prime}=O(sk/\varepsilon^{6})+\text{poly}(k/\varepsilon)$.
@@
    The statements above also hold with $\widetilde{X}_{1}=X_{1}$, but $s^{\prime}=O(sk/\varepsilon^{12})+\text{poly}(k/\varepsilon)$.
@@
As elaborated upon in Section 1.3, the above theorem follows from assigning $\widetilde{X}_{0}$ and $\widetilde{X}_{1}$ to be a careful “mixture” of $X_{0}$ and $X_{1}$. This mixture depends on carefully partitioning the domain space of the random variables using multicalibration. With these random variables, Item 2 follows directly from Item 1 via hybrid argument, so the value of Theorem 1.1 is that it provides not only an upper bound on (in)distinguishability but a matching lower bound, via Item 3. Thinking of $s=n^{\omega(1)},k\leq\mathrm{poly}(n),\varepsilon=1/s$, we see that up to a negligible additive term $(k\varepsilon)$, and a polynomial change in circuit size ($s^{\prime}$ vs. $s$), the computational indistinguishability of $X_{0},X_{1}$ under multiple samples is tightly captured by the information-theoretic indistinguishability of $\widetilde{X}_{0},\widetilde{X}_{1}$ under multiple samples. In particular, $k=\Theta\left(1/d^{2}_{\mathrm{H}}(\widetilde{X}_{0},\widetilde{X}_{1})\right)$ samples are both necessary and sufficient to distinguish $X_{0},X_{1}$ with constant advantage. We can abstract this latter corollary as follows:
@@
    $X_{0}^{\otimes k},X_{1}^{\otimes k}$ are $(1-e^{-k\delta^{2}}-2k\varepsilon)$ distinguishable by circuits of size $s^{\prime}=O(sk/\varepsilon^{6})+\mathrm{poly}(k/\varepsilon)$.
Thus, the number of samples needed to distinguish $X_{0},X_{1}$ with constant advantage is $\Theta(1/\delta^{2})$, where $\delta$ is the pseudo-Hellinger distance. As an immediate consequence, just like the traditional notion of Hellinger distance in the statistical distinguishability setting, our notion of pseudo-Hellinger distance gives an instance-optimal characterization of the computational distinguishability of any pairs of random variables. This is a key benefit of our work over prior works [HR08, Gei22].
@@
    $X_{0}^{\otimes k},\mathcal{U}^{\otimes k}$ are $(1-e^{-\Omega(k\min\{\delta,1\})}-2k\varepsilon)$ distinguishable by circuits of size $s^{\prime}=O(sk/\varepsilon^{12})+\mathrm{poly}(k/\varepsilon)$.
@@ 1.2 Asymptotic Complexity Formulations
We say that ensembles $X,Y$ have computational sample complexity at least $k:\mathbb{N}\rightarrow\mathbb{N}$ for $k(n)=n^{O(1)}$ if $X^{k}\equiv^{\mathrm{comp}}_{1/2}Y^{k}$ (where $X^{k}=\{X_{n}^{k(n)}\}$) [^2] This choice of $1/2$ is arbitrary, and can be amplified through repetition.. We denote this as $X\equiv_{k}^{cS}Y$.
@@
    If $X\equiv_{k}^{cS}Y$, then $X\equiv^{cH}_{\delta}Y$ for some $\delta(n)=O\left(\frac{1}{\sqrt{k(n)}}\right)$.
@@ 1.3 Technical Overview
Our work builds on the recent work of Casacuberta, Dwork, and Vadhan [CDV24]. Driving inspiration from [DLLT23], they showed how the recent notion of multicalibration from the algorithmic fairness literature [HKRR18] leads to simpler and more illuminating proofs of a number of fundamental known results in complexity theory and cryptography, such as the Impagliazzo Hardcore Lemma [Imp95], the Dense Model Theorem [RTTV08], and characterizations of pseudoentropy [VZ12]. We show how multicalibration can be used to derive new results about computational indistinguishability (Theorems 1.1 and 1.5) and in general reduce computational reasoning to information-theoretic reasoning. Specifically, applying the Multicalibration Theorem [HKRR18] to distinguishing problems in a way inspired by [CDV24], we obtain the following. Let $\widetilde{X}_{b|p(\widetilde{X}_{b})=y}$ be the distribution where $x$ is sampled according to $\widetilde{X}_{b}$ conditioned on $p(x)=y$, for some function $p:\mathcal{X}\to[m]$. Let $p(X_{0})$ be the distribution over $[m]$ where for $y\in[m]$, the probability that $p(X_{0})=y$ is exactly $\sum_{x\in p^{-1}(y)}X_{0}(x)$.
@@
    $p$ is computable by circuits of size $O(s/\varepsilon^{6})$.
@@ 1.3.1 Multicalibrated Partitions
We use a partition-based characterization of multicalibration as used in previous works [GKSZ22, GKR+22, DLLT23, CDV24]:
@@
Let $\mathcal{X}$ be a finite domain, for an integer $s\in\mathbb{Z}^{+}$, let $\mathcal{F}^{(s)}$ be the class of functions $f:\mathcal{X}\rightarrow\{0,1\}$ computable by size $s$ circuits, let $g:\mathcal{X}\rightarrow[0,1]$ be an arbitrary function, $\mathcal{D}$ be a probability distribution over $\mathcal{X}$, and $\varepsilon,\gamma>0$. Then, there exists a $(\mathcal{F}_{s},\varepsilon,\gamma)$-approximately multicalibrated partition $\mathcal{P}$ of $\mathcal{X}$ for $g$ on $\mathcal{D}$ such that $\mathcal{P}\in\mathcal{F}^{(W,k)}$, where $W=O(s/\varepsilon^{6})+O(1/(\varepsilon^{12}\gamma)\cdot\log(|\mathcal{X}|/\varepsilon))$ and $k=O(1/\varepsilon)$.
@@
Recently, this perspective on multicalibration has proven to be quite fruitful, with applications to machine learning [GKR+22], graph theory [DLLT23], and complexity theory and cryptography [DLLT23, CDV24]. Philosophically, the work of [CDV24] showed how to reprove (and strengthen) theorems of average-case hardness and indistinguishability using multicalibration (specifically, the Impagliazzo Hardcore Lemma [Imp95], the Dense Model Theorem [RTTV08], and characterizations of pseudoentropy [VZ12]). One consequence of their work was a “loose” intuition that multicalibration translates questions about computational indistinguishability into questions about statistical indistinguishability. Our goal is to show this translation more clearly and generally, and underline directly how multicalibration reduces an understanding of computational indistinguishability to an understanding of statistical indistinguishability (as captured by Theorem 1.1).
@@ 1.3.2 Invoking Multicalibration
With this definition, we can provide an informal overview of the proof of Theorem 1.12. We start with item (b), as its proof is essentially definitional. Recall that this item states that $p(X_{0})$ and $p(\widetilde{X}_{0})$ are identically distributed (and analogously for $X_{1}$). This follows because in the construction of the random variable $\widetilde{X}_{0}$, we sample an element from part $P_{i}$ with probability $\Pr[X_{0}\in P_{i}]$, and thus necessarily, the probability mass placed on each part $P_{i}$ is identical between the two distributions. Next, we can also see that part (c) is definitional. In the construction of $\widetilde{X}_{0}$ and $\widetilde{X}_{1}$, conditioned on being in the piece $P_{i}$, each of the marginal distributions is exactly $\mathcal{D}_{P_{i}}$, and thus the marginal distributions match. Next, to conclude item (d) of Theorem 1.12, recall that the partition function $p$ is exactly returned by the Multicalibration Theorem. Here, we can take advantage of the fact that multicalibrated partitions are efficiently computable given oracle access to the underlying set of test functions (in this case, size $\leq s$ circuits). By replacing these oracles with the size $\leq s$ circuits, this yields an overall complexity of $O(s/\varepsilon^{2})$ for the circuit size required to compute $p$. Finally, it remains to prove item (a) of Theorem 1.12. Because the partition $\mathcal{P}$ that is returned is multicalibrated with respect to our function $g$, by applying Theorem 2.15 of [CDV24] and following the approach of the DMT++++ proof of [CDV24], we can show that for every part $P_{i}\in\mathcal{P}$ and every circuit of size $\leq s$, the distributions $X_{0}|_{P_{i}}$ and $\widetilde{X}_{0}|_{P_{i}}$ are close to indistinguishable (and likewise for $X_{1}$). By combining this “local” indistinguishability across all parts of the partition, this then yields the overall indistinguishability between $\widetilde{X}_{0},\widetilde{X}_{1}$.
@@
It is worth comparing our use of the Multicalibration Theorem to the use of Impagliazzo’s Hardcore Theorem [Imp95, Hol05] in some past works on computational indistinguishability and pseudorandomness [HR08, MT10]. Applied to the same function $g$ above, the Hardcore Theorem says that if $X_{0},X_{1}$ are $(s^{\prime},1-\delta)$ -indistinguishable, for $s^{\prime}=O(s\cdot\mathrm{poly}(1/\varepsilon,1/\delta))$, then there is a subset $H\subseteq X$ such that $\Pr_{x\in X_{0}}[x\in H]=\Pr_{x\in X_{1}}[x\in H]=\delta$ such that $X_{0}|_{H}$ and $X_{1}|_{H}$ are $(s,\varepsilon)$-indistinguishable. Replacing $X_{0}|_{H}$ and $X_{1}|_{H}$ with their average as above, we obtain $\widetilde{X}_{0},\widetilde{X}_{1}$ that are $(s,\varepsilon)$-indistinguishable from $X_{0},X_{1}$ respectively, and such that $d_{\mathrm{TV}}(\widetilde{X}_{0},\widetilde{X}_{1})\leq 1-\delta.$ Thus, the Hardcore Lemma tightly captures the single-sample computational indistinguishability of $X_{0}$ and $X_{1}$ by the single-sample information-theoretic indistinguishability of $\widetilde{X}_{0}$ and $\widetilde{X}_{1}$. But it does not seem to capture enough information to obtain an instance-optimal characterization of the multiple-sample indistinguishability, as we do. Concretely, the Hardcore Lemma requires assuming some initial average-case hardness of $g$ (which comes from the weak indistinguishability of $X_{0}$ and $X_{1}$) and only gives us indistinguishability from an information-theoretic counterpart on a small subset $H\subseteq\mathcal{X}$, which is not guaranteed to be efficiently recognizable. In contrast, the partition given by the Multicalibration Theorem covers the entire domain $\mathcal{X}$ with efficiently recognizable sets, and this is crucial for our instance-optimal characterization.
@@ 2.1 Distinguishability
$d_{H}^{2}\left(X,\mathcal{U}\right)=1-\sqrt{\frac{2^{\mathrm{H}_{1/2}\left(X\right)}}{|\mathcal{X}|}}.$
@@
$d_{H}^{2}\left(X,\mathcal{U}\right)=\frac{1}{2}\sum_{i\in\mathcal{X}}\left(\sqrt{\mathbb{P}[X=i]}-\frac{1}{\sqrt{|\mathcal{X}|}}\right)^{2}=\frac{1}{2}\left(1+1-2\sqrt{\frac{\mathbb{P}[X=i]}{|\mathcal{X}|}}\right)$
$1-\frac{1}{\sqrt{|\mathcal{X}|}}\sum_{i\in\mathcal{X}}\sqrt{\mathbb{P}[X=i]}=1-\sqrt{\frac{2^{\mathrm{H}_{1/2}\left(X\right)}}{|\mathcal{X}|}}.$
@@ 2.2 Multicalibration
Let $\mathcal{X}$ be a finite domain, $\mathcal{F}$ a class of functions, with $f\in\mathcal{F}:f:\mathcal{X}\rightarrow[0,1]$, $g:\mathcal{X}\rightarrow[0,1]$ an arbitrary function, $\mathcal{D}$ a probability distribution over $\mathcal{X}$, and $\varepsilon,\gamma>0$. Then, there exists a $(\mathcal{F},\varepsilon,\gamma)$-approximately multicalibrated partition $\mathcal{P}$ of $\mathcal{X}$ for $g$ on $\mathcal{D}$ such that $\mathcal{P}\in\mathcal{F}_{q,t,k}$ for $t=O(1/(\varepsilon^{12}\gamma)\cdot\log(|\mathcal{X}|/\varepsilon))$, $q=O(1/\varepsilon^{6})$, and $k=O(1/\varepsilon)$.
@@
    $p$ is computable by functions in $\mathcal{F}_{O(1/\varepsilon^{6}),O(1/(\varepsilon^{12})\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))}$.